# Primal-Dual iLQR

### v4

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual Newton-KKT system. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6]. As our algorithm avoids sequential rollouts of the nonlinear dynamics, it can be combined with [17] to run in $O(\log(N))$ parallel time per iteration (where $N$ is the number of stages), as well as $O(1)$ parallel time per line search iteration. Therefore, this paper provides a practical, theoretically sound, and highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
@@ II-C Contributions
Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n))$ parallel time complexity, using a minor variation of [18], and that the remaining parts of our method have constant parallel time complexity.
@@ III-C Merit Functions
Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above. Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
@@ X Parallel Dual LQR
XI Algorithm Overview
In this section, we provide some pseudocode for the sequential version of the proposed algorithm. For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
def build_lqr(problem, x, u, v):
    n = problem.num_states()
    m = problem.num_controls()
    N = problem.num_stages()
    for i in range(N):
        cost_grad = problem.cost(i)
            .gradient()
            .eval(x[:,i], u[:,i])
        dyn_grad = problem.dynamics(i)
            .gradient()
            .eval(x[:,i], u[:,i])
        A_i = dyn_grad.T[:,:n]
        B_i = dyn_grad.T[:,n:]
        if i == 0:
            d_i = problem.initial_state()
                - x[:,0]
        else:
            y = problem.dynamics(i-1)
                .eval(x[:,i-1], u[:,i-1])
            d_i =  y - x[:,i]
        l_i = cost_grad +
            dyn_grad * costates[:,i+1]
            - costates[:,i]
        cost_hess = problem.cost(i)
            .hessian()
            .eval(x[:,i], u[:,i])
        dyn_hess = problem.dynamics(i)
            .hessians()
            .eval(x[:,i], u[:,i])
        Q_i = cost_hess
            + dyn_hess * costates[:,i+1]
    l_N = problem.cost(N).gradient()
        .eval(x[:,N]) - costates[:,N]
    Q_N = problem.cost(N).hessian()
        .eval(x[:,N])
    return LQR(A_*, B_*, d_*, Q_*, l_*)
def lqr_dual_solve(P, p, dx):
    Np1 = dx.shape[1]
    return [P[i] * dx[:, i] + p[i]
        for i in range(Np1)]
def merit_rho(d, dv):
    if d == 0:
        return 0.01
    return 2 * dv.norm() / d.norm()
def eval_merit(problem, x, u, v, rho):
    g = problem.total_cost().eval(x, u)
    d = problem.total_constraints()
        .eval(x, u)
    return g + v.T * d
        + 0.5 * rho * d.squared_norm()
def slope(lqr, dx, du, d, dv, rho):
    N = lqr.num_stages()
    slope = -rho * d.squared_norm()
    for i in range(N):
        dxu = stack(dx[:,i], du[:,i])
        slope += lqr.q(i).T * dxu
        slope += dv[:, i].T * lqr.d(i)
    slope += lqr.q(N).x().T * dx[:, N]
    slope += dv[:, N].T * lqr.d(N)
    return slope
def line_search(problem, lqr, x, u, v,
                dx, du, dv, rho):
    alpha = 1
    m_o = eval_merit(
        problem, x, u, v, rho)
    slope = slope(lqr, dx, d, dv, rho)
    while alpha > kMinStepSize:
        x_n = x + alpha * dx
        u_n = u + alpha * du
        v_n = v + alpha * dv
        m_n = eval_merit(
            problem, x_n, u_n, v_n, rho)
        if m_n < m_o +
            kArmijo * alpha * slope:
            return alpha
    raise LineSearchFailure()
def primal_dual_ilqr(problem, states_ws,
        controls_ws, costates_ws):
    x = states_ws
    u = controls_ws
    v = costates_ws
    for it in range(kMaxIterations):
        lqr = build_lqr(problem, x, u)
        dx, du, P, p = lqr.Solve()
        dv = lqr_dual_solve(P, p, dx)
        if dx == 0 and du == 0:
            if dv == 0:
                return x, u, v
            else:
                v += dv
                continue
        rho = merit_rho(lqr.d(), dv)
        alpha = line_search(
            problem, lqr, x, u,
            v, dx, du, dv, rho)
        x += alpha * dx
        u += alpha * du
        v += alpha * dv
    return x, u, v
XII Benchmarks
@@
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. This seems to be the problem used in [16]. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step. A JAX implementation of our algorithm takes 46ms to finish solving this problem (on the CPU of an M2 MacBook Air). Below, you can find a visual description of the problem and solution, as well as solver logs.
@@
XIII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners. Moreover, it inherits all the desirable properties that NPSQP has been shown to have, such as not impeding local superlinear convergence, global convergence, and the lack of need for second order corrections. None of acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6] combine all of these properties.
The proposed algorithm has sequential time complexity $O(N(n+m)^{3})$, where $N,n,m$ are respectively the number of stages, states, and controls of the given problem, and parallel time complexity $O(\log(m)+\log(N)\log(n)$. Note that a naïf application of the NPSQP algorithm would have sequential time complexity $O(N^{3}(n+m)^{3})$ and parallel time complexity $O(\log(N)+\log(n+m))$ (although requiring a much larger number of workers), as the Newton-KKT systems have dimension $(N+1)(n+m)$. This is due to the fact that inverting a matrix of dimension $d$ takes $O(d^{3})$ sequential time and $O(\log(d))$ parallel time.
This framework can easily be extended to also handle constrained problems. Ideally, an augmented Lagrangian [5] method would be used to handle general equality constraints and an interior point [12] method would be used to handle general inequality constraints. It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian method. Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton. Moreover, as it never performs dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.

### v5

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual Newton-KKT system. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6]. The solutions of the LQR-shaped subproblems posed by our algorithm can be be parallelized to run in time logarithmic in the number of stages, states, and controls. Moreover, as our method avoids sequential rollouts of the nonlinear dynamics, it can run in $O(1)$ parallel time per line search iteration. Therefore, this paper provides a practical, theoretically sound, and highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems. An open-source JAX implementation of this algorithm can be found on GitHub ($joaospinto/primal\textunderscore dual\textunderscore ilqr$).
@@ II-C Contributions
Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n)+\log(m))$ parallel time complexity, using a minor variation of [17], and that the remaining parts of our method have constant parallel time complexity.
@@ III-C Merit Functions
Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above; moreover, $D(\cdot;\cdot)$ is used to represent the directional derivative operator. Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
@@ X Parallel Dual LQR
XI Benchmarks
@@
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from GitHub ($google/trajax$) and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. This seems to be the problem used in [16]. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step. A JAX implementation of our algorithm takes 46ms to finish solving this problem (on the CPU of an M2 MacBook Air). Below, you can find a visual description of the problem and solution, as well as solver logs.
@@
XII Conclusion
This paper introduces a new algorithm for solving unconstrained discrete-time optimal control problems, called Primal-Dual iLQR.
This framework can easily be extended to handle constrained problems. Ideally, an interior point [12] method would be used to handle general inequality constraints. It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian [5] method, although local superlinear convergence would be lost. In both cases, the LQR shape of the posed subproblems can be preserved. Preserving superlinar local convergence in the presence of general equality constraints would require employing a linear equality-constrained LQR subproblem solver. Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method (including adding substantially slower iterative subproblem solves).
Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton, as it treats both state and control trajectories as free variables in the optimization problem (without incurring extra computational costs); this allows them to be independently seeded, even in dynamically infeasible ways.
The proposed algorithm is globally convergent and does not impede superlinear local convergence, without requiring second order corrections to be applied.
Finally, solving the subproblems posed by our method can be parallelized to run in time logarithmic in the number of stages, states, and controls. Moreover, as our method never performs nonlinear dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.